1. Consider three acids: HA with pKa=3, HB with pKa=6, and HC with pKa=9.

(a) Which acid is most appropriate for preparing a buffer at pH=4? Explain briefly.

(b) You are provided with a 0.1 M solution of HB, a 0.1 M solution of B-, a 1 M solution of HCl, and a 1 M solution of NaOH. Describe (without the amounts) the three ways you could prepare a buffer at pH=6 using mixtures of these solutions?

(c) Consider the following: you've prepared a buffer, but the pH is a little higher than you expected. How can you correct this buffer?
Answer #1

Answer (a) :- Acid HA with pKa=3 is most appropriate.

explanation:

The very basic and simple approach, in preaparing a buffer of a particular pH , is that pH of the solution should be in between the buffering range of respective buffer solution.

Now it comes to a question what is buffering range ?

Simply Buffering range is a pH range in which the solution can resist drastic changes in pH when when small amount of an acid or a base is added.

And this Buffering range of the buffer solution is equal to

pKa  +_ 1 or (pKa-1 to pKa+1)

so As per the Q(a) part , if we have to preappr a buffer with pH =4 , then we will check the different acids pKa value in buffering range and the very at the pka wherein pH=4 is , will be our Answer

So for HA, pKa+1 = 3 + 1 = 4  and pKa-1 = 3-1 = 2 ie (2,4)

for HB , pKa+1 = 6+1 = 7 and pKa-1= 6-1 = 5 ie (5, 7)

for HC similarly it is ( 8, 10)

it is clear 4 comes in (2,4) corresponding to pKa =3 of acid HA

so HA is most appreciate.

___________________________________________________________________

Answer (c):-

In general we can prepare a buffer in following two ways

1. Buffer solution with an Acid and it's Conjugate Base

if buffer is prepared by the very way , we can adjust its pH by adding a strong base like NaOH.

2. Buffer solution with Base and conjugate Acid

If buffer is prepare by the above manner its pH can be ajusted by the strong acid like HCl.

therefore if higher pH is to be reduced then any acid as per the compostion of buffer is added.